### v3

@@ Results
The data assimilation for real observational data is more challenging than simulation experiments and has considerable practical application potential. To validate the performance in real-world scenarios, we apply FengWu-Adas to the assimilation for observations of Global Data Assimilation System (GDAS). The values of state variables are interpolated onto the grid points to generate the observations, and the confidence matrix is produced according to the availability and quality of the observations. Observing that satellite-derived data often have larger systematic errors, we conducted two sets of experiments to assimilate all data and non-satellite data of GDAS. Fig. 5 illustrates the variations in RMSE and Bias of the analyses for main variables over the year, where the black crosses indicate that the observational data at the corresponding time are missing. When FengWu-Adas is applied to the assimilation for real observational data, the system can still operate stably for long term, and the RMSE of some variables such as specific humidity can reach a level comparable to that of simulation experiments and the IFS (Extended Data Fig. 11). The results show that although satellite-derived data are less accurate, they can provide more information to improve the quality of the analysis, and it is expected that neural networks are robust to noise. This holds true for most variables except for the pressure level of 50hPa, which may be caused by the small number of high-altitude observational data. What’s more, Adas is robust to handle the lack of observations at a certain time. When the input observations is an all-zeros matrix, it can perform a function like post-processing to maintain the operation and produce the analysis and forecasts.
The average RMSE skill for weather forecasting based on the analyses over ten days is shown in Fig. 6, and the intersections with dashed line correspond to the skillful forecast lead times with ACC=0.6 for z500. The analyses generated after assimilating all data and non-satellite data are used as the initial states for weather forecasting, and the skillful forecast lead times reach 8.75 days and 8.25 days, respectively. The results reflect the forecasting skill of our system after assimilating real observational data, and we have achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time exceeding that of the IFS (8.5 days [43]) for the first time. The GDAS experiments demonstrate the stability and considerable potential for practical applications of our system.
@@ Discussion
There are two main reasons that make assimilating real observational data more challenging than simulated observations. On one hand, the distribution patterns of observational data are different between these scenarios. In simulation experiments, observations generated through random sampling or other methods are approximately distributed uniformly, which is an ideal distribution pattern that yields favorable performance. However, observational data often exhibit a severely non-uniform distribution due to various geographical and human factors in real-world scenarios. For instance, there is a scarcity of observations in polar regions in most cases, which can significantly impact the regional analysis. Such impacts may be propagated through the system’s operation, potentially leading to error accumulation or even collapse of the system. On the other hand, there exists an inherent inconsistency between observational values in the observational space and grid-point values in the grid space. Observational values represent the true and objective reflection of individual states at specific points in space, while grid-point values represent the overall states of a region. Therefore, observational values tend to exhibit larger fluctuations, while grid-point values are more stable and smooth theoretically. This inconsistency can be alleviated as the resolution of grid space increases and the quantity of observational data grows, but it remains a problem that cannot be ignored at present. Traditional data assimilation algorithms address this inconsistency by aligning the grid space to the observational space using observation operator. In contrast, our method requires an operator for the inverse mapping, which may be achieved by adding a pre-processing module instead of simple interpolation and further enhance the performance of our method. In summary, there exists a natural gap between simulated experiments and real-world systems. While simulation experiments are necessary for academic research and algorithm improvement, the ability to be extended to real-world scenarios is of paramount importance for practical applications. We have achieved stable long-term assimilation of real observational data for the first time, which can help forecasting models get rid of the dilemma of relying on the initial states provided by inefficient traditional data assimilation components, and promote the rapid development of research in related fields.
The future work will mainly focus on a more universal and high-performance framework. For example, techniques based on continuous space modeling [44] may provide a train of thought for better solution, which can work directly with the off-grid real observations without pre-processing or observation operator that may introduce additional errors. Nowadays, high-quality data has always been essential for the development of AI and many other technologies, and the reanalysis datasets still hold an irreplaceable significance. Therefore, developing physical models and traditional algorithms remains crucial. The purpose of our study is not to replace the traditional NWP systems but to demonstrate AI methods’ potential in tackling real-world challenges. Both physical techniques and AI methods have their own strengths and limitations, and they should be treated equally and considered as powerful tools for benefiting science and society. We believe that this is a meaningful exploration and look forward to the deployment and implementation of AI methods in operational systems in the future.
@@ Methods
ERA5 is used as the ground truth and source of the simulated observations in our experiments. ERA5 is a global atmospheric reanalysis archive containing hourly weather variables such as temperature, geopotential, wind speed, humidity, etc. A subset of ECMWF’s ERA5 dataset for 40 years, from 1979 to 2018, is chosen to train and evaluate the model. We choose to conduct experiments on a total of 69 variables at a resolution of 0.25° ($721\times 1440$ grid points), including five upper-air variables with 13 pressure levels (i.e., 50hPa, 100hPa, 150hPa, 200hPa, 250hPa, 300hPa, 400hPa, 500hPa, 600hPa, 700hPa, 850hPa, 925hPa, and 1000hPa), and four surface variables. Specifically, the upper-air variables are geopotential (z), temperature (t), specific humidity (q), zonal component of wind (u) and meridional component of wind (v), whose 13 sub-variables at different vertical level are presented by abbreviating their short name and pressure levels (e.g., z500 denotes the geopotential at a pressure level of 500 hPa), and the surface variables are 10-meter zonal component of wind (u10), 10-meter meridional component of wind (v10), 2-meter temperature (t2m) and mean sea level pressure (msl). Following a common protocol in the simulation experiments, the data from 1979-2015 are used for training, 2016-2017 for validation, and 2018 for testing.
The real observational data is parsed from the prepbufr files of GDAS, which is archived by National Centers for Environmental Information (NCEI). GDAS provides various types of observational data starting from 2012-02-13 00:00UTC, including multi-source data such as surface land reports, surface marine reports, satellite-derived wind reports, etc. Each datum contains the longitude and latitude coordinates, pressure level, time, and the value and quality marker of state variables. Regarding the large number of missing GDAS observations after 2018, we use the data from 2017 for testing and others for training to demonstrate the quality of the analysis for the entire year.
@@
The latitude-weighted root mean square error (RMSE) and the latitude-weighted anomaly correlation coefficient (ACC) are statistical metrics widely used in geospatial analysis and atmospheric science, which are both computed against ground truth data. And latitude weighting is a common strategy to account for the varying area represented by different latitudes on a spherical Earth. RMSE evaluates the accuracy of a model’s forecasts or estimates of variables across different latitudes, and ACC reflects the model’s ability to capture anomalies (departures from the long-term averaged climatology). Regarding the bias drift of AI models, the latitude-weighted mean error (Bias) is also used to assess the bias of estimates against the ground truth [17], and the Bias of an ideal system should be close to zero. The definitions of the metrics are as below.
RMSE
@@
where $T$ is the total number of test time slots.
ACC
Given the forecast $\hat{x}_{h,w,c}^{i+\tau}$ and its ground truth $x_{h,w,c}^{i+\tau}$, and the climatological mean over the day-of-year $C_{h,w,c}^{i+\tau}$ for the $c$-th channel with lead time $\tau$, the ACC is defined as
$\operatorname{ACC}(c,\tau)=\frac{1}{T}\sum_{i=1}^{T}\frac{\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})(\hat{x}_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})}{\sqrt{\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})^{2}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(\hat{x}_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})^{2}}}$
where $C_{h,w,c}^{i+\tau}$ is averaged from 1993 to 2016 with ERA5 data [13, 14].
Bias
Given the estimate $\hat{x}_{h,w,c}^{t}$ and its ground truth $x_{h,w,c}^{t}$ for the $c$-th channel at time $t$, the Bias is defined as
$\operatorname{Bias}(c,t)=\frac{1}{H_{0}\cdot W_{0}}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{t}-\hat{x}_{h,w,c}^{t})$
@@
The patch embedding and patch recovery are implemented through standard convolution and transposed convolution, respectively. The patch size of Adas in our experiments is set to 6 and the number of vertical layers remain unchanged. This means a patch has $1\times 6\times 6$ pixels for 3D upper-air variables and $6\times 6$ pixels for 2D surface variables. The kernel sizes and strides of 3D and 2D convolution are the same as the patch size, and the two parts are then concatenated together in the vertical direction. Therefore, in our experiments, $H=H_{0}/6$, $W=W_{0}/6$, $D=14$, and the dimension $C$ is 192. The parameters of patch embedding and patch recovery are not shared for each input.
@@
In order to capture multi-scale meteorological features, the patch merging and patch expanding modules are used for down-sampling and up-sampling (Fig. 2d). The patch merging firstly rearrange pixels to reduce the horizontal dimensions by half, while the number of channels increases. After the regularization layer, the number of channels is then transformed by a linear layer. Therefore, the patch merging achieves halving of the horizontal dimensions and doubling of the channel dimension. The patch expanding module performs the opposite operation and makes the encoder form a symmetrical Unet [45] structure. Similarly, the number of vertical layers is not changed in the process of patch merging and patch expanding.
@@
Gated convolution is a convoluional operation based on soft-gating, which employs continuous masks ranging between 0 and 1 to represent the degree of validity for each pixel. Originally, the gating mask in gated convolution was learned automatically from data at each layer, and the output features were weighted by it. In Adas, the confidence matrix of our method can serve as the gating mask. By using the confidence matrix as input for each layer and updating it explicitly, we propose an improved form of gated convolution to provide a more complete representation for sparse observations. The guidance of the confidence matrix can help the network extract features more effectively and capture the interactions between the background and observations more efficiently in the gated cross-attention module. Specifically, each value in the confidence matrix signifies the degree of confidence associated with corresponding observation values, and the confidence matrix is dynamically updated to guide the varying confidence levels of different layers. Let the subscript $i$ and $o$ represent the input and output of each layer, respectively, then the gated convolution and the update rule of the confidence matrix can be formulated as
@@
where $\odot$ represents element-wise multiplication. The $Sigmoid(\cdot)$ activation ensures that the values of the confidence matrix are between 0 and 1, and the activation funtion for features can be arbitrary. We use the $SiLU(\cdot)$ activation here, which has been proven to be a better choice than $ReLU(\cdot)$ [46, 47]. The gated convolution module can provide a more complete representation for observations and improve the quality of feature fusion.
@@
The cross-attention mechanism is introduced to capture the interactions between the background and observations, but treating observations at different locations with different confidence levels equally is clearly unreasonable. Especially for observations with low confidence levels, directly calculating cross-attention may have negative guidance on the background. To address this issue, we propose the gated cross-attention module, which is also guided by the confidence matrix. Specifically, when the observations are used as the condition of cross-attention, the background is aligned with them at first, ensuring that observations influence the background in proportion to the confidence levels. After calculating the cross-attention, the remaining proportion of the background that does not participate in the operation will be added back. Conversely, the background fully influences observations since all values in the background have relative high confidence levels generally, which can correct the observations with low confidence levels. Therefore, the confidence matrix needs to be updated accordingly, which is also implemented through a convolutional layer with $Sigmoid(\cdot)$ activation (Fig. 2c). The process of gated cross-attention can be expressed as
@@
The mean absolute error (MAE) loss, also known as L1 loss, is employed to supervise the training of the neural network. In multivariate optimization, an issue of imbalanced optimization arises when there are significant differences in the magnitudes of losses for different variables. Therefore, it is necessary to weigh the losses for different variables. Instead of setting the weights manually, the losses for each variable (i.e., each channel) are automatically weighted to have the same magnitude in our training. The models in simulation experiments are trained for 50 epochs using the AdamW optimizer [48] and OneCycleLR scheduler [49]. The learning rate starts from 1e-6, warms to a maximum value of 1e-4 for ten epochs, and then decays gradually to 1e-10. To ensure adequate training with available GDAS observational data, the models in GDAS experiments are fine-tuned for 50 epochs based on the transfer learning strategy.
@@
Compared to traditional methods, the reduction of computational consumption and execution time is a significant advantage of neural networks. In our system, Adas requires only about 0.6 seconds for the inference of data assimilation on an NVIDIA Tesla-A100 GPU, and FengWu needs about 1.5 seconds for inferring a single-step forecast. That is to say, our system can complete a periodic iteration in seconds once the observational data are available, which achieves significant efficiency improvement (about 1000-10000 times faster) compared to traditional data assimilation algorithms [50].
[…]

### v4

@@ Results
Data assimilation for real observational data is more challenging than simulation experiments and has considerable practical application potential. To validate the performance in real-world scenarios, we apply FengWu-Adas to the assimilation for conventional observations from the Global Data Assimilation System (GDAS). The main data types and information of GDAS observations we use are given in Table 1. The values of geophysical variables are processed into grid data through nearest neighbor interpolation to generate the observation fields, and the confidence matrix is produced according to the availability and quality of the observations (details in Methods). Figure 4a and Figure 4b illustrate the RMSE variations of the analyses for $z500$ and $t850$ variables over the year in the cyclic forecast-assimilation experiment, which are calculated against ERA5 and the 200 reserved GDAS observation columns, respectively. When FengWu-Adas is applied to the assimilation for real observational data, the system can still operate stably for long term. Compared with the baseline of 3D variational (3DVar) data assimilation, Adas has lower analysis error and stronger stability. It is worth noting that the 3DVar we reproduced may not be optimal, and it may achieve improved performance with careful tuning. It is encouraging that the analysis quality of Adas evaluated at the stations shows a comparable level to IFS-Analysis and ERA5, although Adas only assimilates conventional observations. Furthermore, we analyze the statistical characteristics of the RMSE across the evaluation stations at all times and draw the box plot in Figure 4c. Whether considering the median, interquartile range (IQR) or outliers, Adas shows reasonable results and is not significantly worse than IFS-Analysis and ERA5. Figure 4d shows the scatter plot and corresponding fits between the observation and analysis values across the evaluation stations at all times, with the coefficient of determination $R^{2}$ and the slope of the fitted line. The analysis values of Adas fit the observations well, and its $R^{2}$ is only slightly lower than that of IFS-Analysis and ERA5. Figure 5a and Figure 5b visualize the fields of $z500$ and $t850$ variables and their error distributions with ERA5 during data assimilation, respectively, where the background uses the $96h$ lead forecast based on ERA5 and the date-time is randomly selected at 2017-01-27 00:00 UTC. Based on the information provided by observations, Adas can effectively reduce the background error, and the analysis in areas with dense observations will be more accurate. Figure 5c visualizes the RMSE distribution evaluated on the reserved stations at the same date-time, they all show similar errors since the learning target of Adas is ERA5.
Table 1: Main data information of GDAS observations used in FengWu-Adas system. GDAS observational data consists of many messages with different types, each of which contains many observation columns with different height levels. The conventional types of observations in GDAS are used for data assimilation and evaluation in our experiments, covering the main geophysical variables such as temperature, humidity and wind speed. The table shows the number of messages and observation columns of each data type at 2017-01-01 00:00 UTC. Upper-air observations are mainly provided by ADPUPA and AIRCFT, and each observation column of ADPUPA usually contains 1 to more than 100 unfixed height levels. Surface observations of land and ocean are mainly provided ADPSFC and SFCSHP, respectively.
Type Number of Number of Description
Messages Obs. Columns
ADPUPA 338 1403 Upper-Air (RAOB, PIBAL, RECCO,
DROPS) Reports
AIRCFT 21 2255 AIREP/PIREP, AMDAR(ASDAR/ACARS),
E-ADAS(AMDAR BUFR) ACFT
ADPSFC 1117 122093 Surface Land (SYNOP, METAR) Reports
SFCSHP 150 20268 Surface Marine (SHIP, BUOY,
C-MAN Platform) Reports
Figure 4: The performance of FengWu-Adas to produce analysis with GDAS real observations. a, RMSE variations of the analyses for $z500$ and $t850$ variables evaluated on ERA5 and the comparison with 3DVar algorithm. By alternating between forecasting and data assimilation, the system can also maintain long-term stability in real-world scenarios. b, RMSE variations of the analyses evaluated on 200 reserved GDAS observation columns and comparison with IFS-Analysis and ERA5. The analysis quality of Adas evaluated at the stations shows a comparable level to IFS-Analysis and ERA5, although Adas only assimilates conventional observations. c, The box plot of the RMSE across the evaluation stations at all times. Whether considering the median, IQR or outliers, Adas shows reasonable results and is not significantly worse than IFS-Analysis and ERA5. d, The scatter plot and corresponding fits between the observation and analysis values across the evaluation stations at all times.
Figure 5: Visualization of the fields during data assimilation and the RMSE distribution evaluated on the reserved stations. a, Visualization of $z500$ variable and corresponding error distributions with ERA5 during data assimilation. The date-time is randomly selected at 2017-01-27 00:00 UTC and the background uses the $96h$ forecast based on ERA5. b, Visualization of $t850$ variable and error distributions during data assimilation. c, Visualization of the RMSE distribution at the reserved stations and comparison with IFS-Analysis and ERA5. The date-time is also selected at 2017-01-27 00:00 UTC, and the analysis of Adas shows similar errors to IFS-Analysis and ERA5.
To comprehensively evaluate the end-to-end forecasting skill of FengWu-Adas, we test its performance on ERA5, GDAS, and the Integrated Global Radiosonde Archive (IGRA) datasets and compare it with the IFS-HRES. IGRA covers radiosonde observations from more than 2,800 stations around the world (vandal2024global). The evaluation on GDAS is calculated against 1000 stations with highest quality to avoid the impact of poor observations. Figure 6a shows the annual average RMSE skill for $z500$ and $t850$ variables over 10 days. All evaluations show the same trend, with FengWu-Adas having higher forecast RMSE than IFS-HRES in the early stage and lower than IFS-HRES starting around 7-8 days. Since Adas assimilates far fewer observations than the IFS, resulting in less accurate analysis, FengWu-Adas has a larger RMSE for forecasts at shorter lead times. However, FengWu accumulates errors more slowly for multi-step forecasts, so FengWu-Adas has a smaller RMSE for forecasts at longer lead times. Adas mainly provides the initial state for the forecasting model, which contributes more at shorter lead times, and FengWu contributes more at longer lead times. Figure 6b shows the box plot of the prediction RMSE with different lead times. Similar to the results of average RMSE in Figure 6a, FengWu-Adas performs slightly worse than IFS-HRES at shorter lead times, but performs better at longer lead times, with lower medians and IQRs. The scatter plot and corresponding fits between the observation and prediction values across the evaluation stations are illustrated in Figure 7. The first two rows and last two rows correspond to the $z500$ and $t850$ variables, respectively. The columns have different lead times of 1, 5 and 10 days, and the coefficient of determination $R^{2}$ and the slope of the fitted line are marked in the subplot. The prediction of FengWu-Adas is more consistent with the observations at longer lead times and has a higher $R^{2}$. In addition, we visualize the distribution of forecast RMSE at the GDAS evaluation stations in Figure 8. The initialization date-time is also selected as 2017-01-27 00:00 UTC, which is consistent with Figure 5. The first, second and third columns correspond to the RMSE distributions with 1-day, 5-day and 10-day forecast lead times, respectively. It can be seen that in the forecast results with 10-day lead time, IFS-HRES has a large deviation in North America, while FengWu-Adas has a better forecast performance in this region. In general, FengWu-Adas shows accurate end-to-end global weather forecasts, demonstrating the promising potential of data-driven approaches.
Figure 6: The performance of FengWu-Adas to produce end-to-end global weather forecasts and the comparison with IFS-HRES. a, Annual average RMSE skill for $z500$ and $t850$ variables over 10 days evaluated on ERA5, GDAS and IGRA datasets. All evaluations show the same trend, with FengWu-Adas having higher prediction RMSE than IFS-HRES in the early stage and lower than IFS-HRES starting around 7-8 days. FengWu-Adas shows accurate end-to-end global weather forecasts, demonstrating the promising potential of data-driven approaches. b, The box plot of the prediction RMSE across the GDAS evaluation stations with different lead times. Similar to the results of average RMSE, FengWu-Adas performs slightly worse than IFS-HRES at shorter lead times, but performs better at longer lead times with lower medians and IQRs.
Figure 7: The scatter plot and corresponding fits between the observation and prediction values across the evaluation stations. The first two rows and last two rows correspond to the $z500$ and $t850$ variables, respectively. The columns have different lead times of 1, 5 and 10 days, and the coefficient of determination $R^{2}$ and the slope of the fitted line are marked in the subplot. The prediction of FengWu-Adas is more consistent with the observations at longer lead times and has a higher $R^{2}$.
Figure 8: Visualization of the RMSE comparison of FengWu-Adas and IFS-HRES forecasts for $z500$ and $t850$ variables at the GDAS stations. The initialization date-time is set at 2017-01-27 00:00 UTC, which is consistent with Figure 5, and the RMSE is calculated against GDAS observations. The columns have different lead times of 1, 5, and 10 days, and the first two rows and last two rows correspond to the $z500$ and $t850$ variables, respectively. In the forecast results with 10-day lead time, IFS-HRES has a large deviation in North America, while FengWu-Adas has a better forecast performance in this region.
@@ Discussion
Significant progress has been made in AI-based weather forecasting research on the ideal setting based on ERA5 dataset, which prompts our focus to shift towards end-to-end forecasting based on realistic observations. Some ambitious yet radical endeavors have aimed to discard the background and data assimilation components, striving for direct forecasting from observations as initial conditions (vaughan2024aardvark; mcnally2024data). However, to date, the forecast-assimilation cycle remains the most reliable mainstream paradigm for operational end-to-end weather forecasting. This work primarily concentrates on data assimilation and end-to-end weather forecasting based on conventional observations (e.g., in-situ weather stations, ships and buoys, radiosonde observations), which serve as the most direct and accurate information sources. Nowadays, satellite observations are playing an increasingly vital role in data assimilation. Typically, they do not directly provide observations of geophysical variables but rather measurements such as radiation, brightness temperature, reflectivity, etc., which require additional inversion steps, although neural networks inherently suit this process (xu2024fuxi). Compared with the pre-satellite era, the operational forecasting capabilities of meteorological centers using a complete global observing system including satellites have improved dramatically. Lacking the assimilation for satellite and radar observations, especially satellite data that are important for large-scale processes, constitutes a primary limitation of this study and a key direction for our future exploration. Another limitation lies in that our method is based on grid data of a given resolution, and interpolating the original off-grid data will inevitably introduce additional errors. Recently, techniques based on continuous spatial modeling have offered a promising train of thought to directly handle off-grid data and have achieved impressive results in scaling to arbitrary resolutions (chen2024fnp). Moreover, this method of modeling implicit functional representation can support multi-modal data (e.g., in-situ observations and satellite data) to form a universal framework and can be readily extended to 4D data assimilation. In addition, since our method does not change the forecasting model, the end-to-end forecasts of the system still inevitably inherit its shortcomings, such as the blurred prediction. This problem is hopeful to be solved as the forecasting model improves.
In fact, ERA5 itself integrates various observational data, including conventional observations, satellites, and radars, in the process of being generated. Unlike the optimization process of traditional data assimilation methods, supervised learning makes this information inevitably part of the implicit priors introduced in model training. For the current end-to-end paradigm, the precision of the supervised labels determines the upper limit of the method’s performance, posing a challenge to be addressed in the future. Producing high-quality analysis data and forecasts that may be superior to any existing product should be the ultimate goal of data assimilation and data-driven revolution, which still requires a lot of exploration and efforts. Recent study has shown promising progress and may become the mainstream paradigm for the next stage of the research in the domain of end-to-end weather forecasting (alexe2024graphdop). The reanalysis datasets offer the highest-quality complete labels at present and hold an irreplaceable significance. In fact, high-quality data has always been essential for the development of AI and many other technologies. Therefore, the development of physical models and traditional algorithms remains crucial. The purpose of our study is not to replace the traditional NWP systems but to demonstrate the potential of AI methods in tackling real-world challenges. Both physical techniques and AI methods have their own strengths and limitations and should be treated equally and considered as powerful tools for benefiting science and society. We believe that this is a meaningful exploration and look forward to the deployment and implementation of AI methods in operational systems in the future.
@@ Methods
ERA5 is used as the ground truth and source of the simulated observations in our experiments. ERA5 is a global atmospheric reanalysis archive containing hourly weather variables such as temperature, geopotential, wind speed, humidity, etc. A subset of the ERA5 dataset for 39 years, from 1979 to 2017, is chosen to train and evaluate the model. We choose to conduct experiments on a total of 69 variables at a resolution of 0.25° ($721\times 1440$ grid points), including 5 upper-air variables with 13 pressure levels (i.e., 50hPa, 100hPa, 150hPa, 200hPa, 250hPa, 300hPa, 400hPa, 500hPa, 600hPa, 700hPa, 850hPa, 925hPa and 1000hPa), and 4 surface variables. Specifically, the upper-air variables are geopotential ($z$), temperature ($t$), specific humidity ($q$), zonal component of wind ($u$) and meridional component of wind ($v$), whose 13 sub-variables at different vertical level are presented by abbreviating their short name and pressure levels (e.g., $z500$ denotes the geopotential at a pressure level of 500 hPa), and the surface variables are 10-meter zonal component of wind ($u10$), 10-meter meridional component of wind ($v10$), 2-meter temperature ($t2m$) and mean sea level pressure (mslmsl). We use the data from 1979-2015 for training, 2016 for validation, and 2017 for testing.
The real observational data are parsed from the prepbufr files of GDAS, which are archived by the National Centers for Environmental Information (NCEI). GDAS provides various types of observational data starting from 2012-02-13 00:00UTC, including multi-source data such as aircraft and upper-air reports, surface land/marine reports, scatterometer data, satellite-derived wind reports, etc. The prepbufr files consist of many messages with different types, each of which contains many observation columns with different height levels. The conventional types of observations in GDAS are used for data assimilation and evaluation in our experiments, covering the main geophysical variables such as temperature, humidity, and wind speed. Each observation contains the longitude and latitude coordinates, pressure level, time, and the value and quality marker of the geophysical variables. Before being input into the neural network, the observations are first processed into 0.25° grid data through nearest neighbor interpolation, and the grid values without observations are filled with 0. The confidence matrix is specified based on the quality marker, time offset, and interpolation distance to represent the quality of the observation value, and the confidence at places without observations is set to 0 to indicate filled spurious observation values. Regarding the large number of missing GDAS observations after 2018, data from 2017 are used for testing and others for training.
@@
The latitude-weighted root mean square error (RMSE) is a statistical metric widely used in geospatial analysis and atmospheric science, which evaluates the accuracy of a model’s forecasts or estimates of variables across different latitudes. The latitude weighting is a common strategy to account for the varying area represented by different latitudes on a spherical Earth. The definition of RMSE is as follows.
@@
where $T$ is the total number of test time slots. The RMSE calculated at the stations is not latitude-weighted.
@@
The patch embedding and patch recovery are implemented through standard convolution and transposed convolution, respectively. The patch size of Adas in our experiments is set to 6 and the number of vertical layers remains unchanged. This means that a patch has $1\times 6\times 6$ pixels for 3D upper-air variables and $6\times 6$ pixels for 2D surface variables. The kernel sizes and strides of 3D and 2D convolution are the same as the patch size, and the two parts are then concatenated together in the vertical direction. Therefore, in our experiments, $H=H_{0}/6$, $W=W_{0}/6$, $D=14$, and the dimension $C$ is 192. The parameters of patch embedding and patch recovery are not shared for each input.
@@
In order to capture multi-scale meteorological features, the patch merging and patch expanding modules are used for down-sampling and up-sampling (Figure 2d). Patch merging first rearranges pixels to reduce the horizontal dimensions by half, while the number of channels increases. After the regularization layer, the number of channels is then transformed by a linear layer. Therefore, patch merging achieves halving of the horizontal dimensions and doubling of the channel dimension. The patch expanding module performs the opposite operation and makes the encoder form a symmetric Unet (ronneberger2015u) structure. Similarly, the number of vertical layers is not changed in the process of patch merging and patch expanding.
@@
Gated convolution (yu2019free) is a convoluional operation based on soft-gating, which employs continuous masks ranging between 0 and 1 to represent the degree of validity for each pixel. Originally, the gating mask in gated convolution was learned automatically from the data at each layer, and the output features were weighted by it. In Adas, the confidence matrix of our method can serve as the gating mask. By using the confidence matrix as input for each layer and updating it explicitly, we propose an improved form of gated convolution to provide a more complete representation for sparse observations. The guidance of the confidence matrix can help the network extract features more effectively and capture the interactions between the background and observations more efficiently in the gated cross-attention module. Specifically, each value in the confidence matrix signifies the degree of confidence associated with corresponding observation values, and the confidence matrix is dynamically updated to guide the varying confidence levels of different layers. Let the subscript $i$ and $o$ represent the input and output of each layer, respectively, then the gated convolution and the update rule of the confidence matrix can be formulated as
@@
where $\odot$ represents element-wise multiplication. The $Sigmoid(\cdot)$ activation ensures that the confidence matrix values are between 0 and 1, and the activation function for the features can be arbitrary. We use the $SiLU(\cdot)$ activation here, which has been proven to be a better choice than $ReLU(\cdot)$ (hendrycks2016gaussian). The gated convolution module can provide a more complete representation for observations and improve the quality of feature fusion.
@@
The cross-attention mechanism is introduced to capture the interactions between the background and observations, but treating observations at different locations with different confidence levels equally is clearly unreasonable. Especially for observations with low confidence levels, directly calculating cross-attention may have negative guidance on the background. To address this issue, we propose the gated cross-attention module, which is also guided by the confidence matrix. Specifically, when the observations are used as the condition of cross-attention, the background is aligned with them at first, ensuring that observations influence the background in proportion to the confidence levels. After calculating the cross-attention, the remaining proportion of the background that does not participate in the operation will be added back. In contrast, the background fully influences observations since all values in the background generally have relatively high confidence levels, which can correct the observations with low confidence levels. Therefore, the confidence matrix needs to be updated accordingly, which is also implemented through a convolutional layer with $Sigmoid(\cdot)$ activation (Figure 2c). The process of gated cross-attention can be expressed as
@@
The mean absolute error (MAE) loss, also known as L1 loss, is employed to supervise the training of the neural network. In multi-variate optimization, an issue of imbalanced optimization arises when there are significant differences in the magnitudes of losses for different variables. Therefore, it is necessary to weigh the losses for different variables. Instead of setting the weights manually, the losses for each variable (i.e., each channel) are automatically weighted to have the same magnitude in our training. The models in simulation experiments are trained for 50 epochs using the AdamW optimizer (loshchilov2018fixing) and OneCycleLR scheduler (smith2019super). The learning rate starts from 1e-6, warms to a maximum value of 1e-4 for 10 epochs, and then decays gradually to 1e-10. To ensure adequate training with available GDAS observational data, the models in GDAS experiments are fine-tuned for 50 epochs based on the transfer learning strategy.
@@
Compared to traditional methods, the reduction of computational consumption and execution time is a significant advantage of neural networks. In our system, Adas requires only about 0.6 seconds for the inference of data assimilation on an NVIDIA Tesla-A100 GPU, and FengWu needs about 1.5 seconds for inferring a single-step forecast. This means that our system can complete a periodic iteration in seconds once the observational data are available, which achieves significant efficiency improvement (about 1000-10000 times faster) compared to traditional data assimilation algorithms (zhang2019operational).